OpenStudy (anonymous):

(x - y )(x 2 + xy + y 2 )

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (skullpatrol):

$$\Huge (x - y )(x^2 + xy + y^2 )$$

OpenStudy (skullpatrol):

multiply it out and see what you get :-)

OpenStudy (anonymous):

uh distributive property right? but how would I go by multiplying 2 numbers against 3

OpenStudy (anonymous):

Can be thought of distributing 1 thing to two: set A = (x^2 +xy+y^2) then distrubute to (x-y)*A = (x*A - y*A)

OpenStudy (skullpatrol):